# Distinguishing Orbiting and Infalling Dark Matter Particles with Machine Learning

### v1

@@ 3.2. Accuracy of Density Profiles
Figure 5.— Same as Fig. 4, but for the Planck-like cosmology. Similarly to the fiducial cosmology, the model closely matches the density profiles based on Sparta except for regions with very few particles. Unlike the Bolshoi cosmology simulations, the Planck cosmology simulations only reach a box size of $500\>\rm{Mpc}$, which results in too few halos in the $6.03.0<\nu<$ 6.0 bin.
@@
4. Discussion
4.1. Interpreting the ML Model
@@
Figure 6.— Feature importance for each particle in the model’s classifications, as captured by SHAP values. We take two random samples of 5,000 particles from our testing dataset, one from all particles and one from particles that do not have past snapshot information. We examine how the model weighted each feature’s value in its classification. Negative SHAP values mean that a feature is indicative of infalling particles, with positive values indicating orbiting. The color scale represents the range of values for each feature rather than exact numbers, with red points indicating higher values and blue points indicating lower values for a specific particle’s feature. Grey points indicate that the values for that feature are not present, which happens for particles that were not within the search radius in the past snapshot. The radius of the particle at the current and past snapshot has the largest influence on the model’s classification for both populations, indicated by the larger SHAP values the model assigns this feature. The present radial and tangential velocites have similar shapes for both groups while past velocities only significantly contribute with extreme velocities.
Figure 7.— Comparison of kinetic-energy cuts to Sparta’s classification. We investigate cuts in linear (top-row) and log (bottom-row) scaled phase-space of $\ln(v^{2}/v^{2}_{\rm 200m})$ versus $r/R_{200m}$ from all the WMAP7 simulations. The particles are split by their classification according to Sparta and then further by their sign of their radial velocity. The two kinetic-energy cut methods, described in Section 4.2, are overplotted as green lines (Fast) and cyan bins (Optimized). The kinetic-energy cuts classify particles below their lines as orbiting and those above as infalling. The blue line indicates the furthest radial extent that particles are considered for the simple kinetic-energy cuts’ fit. The bin-by-bin fit of the Optimized method disagrees with the best-fit lines significantly, and naturally leads to a more accurate match to Sparta’s classification.
@@
Overall, the model seems to classify particles without past information as infalling, as one would expect for particles that were outside the search radius at that time. The only exceptions are particles that belong to halos that did not exist 0.75 dynamical times ago or had fast enough infalling velocities to have a pericenter despite the distance. Only a small subset of this population eventually orbits the halo, reinforcing our decision to limit the snapshot separation to 0.75 dynamical times. Increasing the time separation would increase the number of particles in this category, reduce the relevance of past snapshots, and provide less information on a particle’s orbiting classification.
4.2. Comparison to Kinetic Energy Cut
@@
Fig. 7 compares the Fast method (green lines) to the distribution from Sparta in isolated halos, defined as halos whose most massive neighbor within $2R_{\rm 200m}$ is at most 20% of the mass of the central halo. We split the particle population into four samples with positive/negative radial velocities and orbiting/infalling classifications from Sparta. The green lines in Fig. 7 show that the linear kinetic-energy cut separates the bulk of the orbiting and infalling populations as classified by Sparta. The figure shows both linear (top) and logarithmic (bottom) scales, highlighting the bulk of the population and outliers, respectively. Figure 8 additionally shows the fraction of orbiting particles as a function of radius as determined by both algorithms. The Fast and Sparta algorithms are in reasonably good agreement, except for the very inner regions and beyond $2R_{\rm 200m}$, where Fast classifies all particles as infalling.
@@
By contrast, the logarithmic scale for the bottom row of Fig. 7 highlights the tails of the distributions, making it easier to see where the Sparta and Fast classifications disagree. The most important difference is due to a population of particles classified as infalling by Sparta at large radii and with low kinetic energy, i.e the particles that fall below the Fast line in the right-most plot of Fig. 7. These particles “boost” the outskirts of the orbiting profile of halos in the Fast algorithm relative to the profile obtained with Sparta. This is illustrated in Fig. 9, where we compare the orbiting and infall profiles of halos as determined using Fast and Sparta. The truncation of the Sparta orbiting profiles is much more sensitive to peak height than the truncation observed using the Fast algorithm. This type of sensitivity is expected: the outer profile is known to be sensitive to accretion rate (Diemer & Kravtsov, 2014), which increases with peak height, suggesting the outer profiles in Sparta are better physically motivated than those obtained using the Fast algorithm.
However, it is not immediately clear whether this is a limitation of the Fast algorithm in particular or kinetic-energy cuts in general. We have thus also tested the Optimized method, which corresponds to finding a radius-dependent kinetic energy cut that minimizes the differences with Sparta’s classification. Fig. 7 shows that the resulting cuts (in cyan) cannot be adequately described as a line. Naturally, the Optimized classification better matches Sparta, but the right columns of Fig. 9 demonstrate that differences near the truncation of the orbiting profiles remain. These differences are fundamentally due to the fact that distant, low-energy infalling particles have phase-space positions indistinguishable from orbiting particles near their apocenter, leading phase-space based classifiers to label these particles as orbiting. However, when adding information from a snapshot that is 0.75 dynamical times earlier, it becomes trivial to separate these particles: infalling particles used to be much further away from the halo, while orbiting particles in the halo outskirts used to be near the halo center at that time. We have explicitly verified this is the case, explaining why our machine-learning method can adequately distinguish between orbiting and infalling particles in the halo outskirts.
@@ Acknowledgments
This research was supported in part by the National Science Foundation under Grant numbers 2206690. BD is grateful to the Sloan Foundation for their financial support. E.S. and E.R. were supported by NSF grant 3047980. E.R. is also funded by DOE grant DE-SC0009913. The Erebos simulations were run on the Midway computing cluster provided by the University of Chicago Research Computing Center, and our analysis was performed on the Zaratan cluster at the University of Maryland. This research extensively used the python packages Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), Matplotlib (Hunter, 2007), Pygadgetreader (Thompson, 2014), XGBoost (Chen & Guestrin, 2016), SHAP (Lundberg et al., 2019), DASK (Dask Development Team, 2016), and Colossus (Diemer, 2018).
@@
Our code is publicly available at github.com/ZevVladimir/ATHENA, including a description of how to create the datasets needed and how to train additional models. We also provide the json of the XGBoost model used and configuration parameters. The full particle data for the Erebos $N$-body simulations are too large to be permanently hosted online, but they are available upon request.

### v2

@@ III.2. Accuracy of Density Profiles
Figure 5.— Same as Fig. 4, but for the Planck-like cosmology. Similarly to the fiducial cosmology, the model closely matches the density profiles based on Sparta except for regions with very few particles. Unlike the Bolshoi cosmology simulations, the Planck cosmology simulations only reach a box size of $500\>\rm{Mpc}$, which results in too few halos in the $3.0<\nu<6.0$ bin.
@@
IV. Discussion
IV.1. Interpreting the ML Model
@@
Figure 6.— Feature importance for each particle in the model’s classifications, as captured by SHAP values. We take two random samples of 5,000 particles from our testing dataset, one from all particles and one from particles that do not have past snapshot information. We examine how the model weighted each feature’s value in its classification. Negative SHAP values mean that a feature is indicative of infalling particles, with positive values indicating orbiting. The color scale represents the range of values for each feature rather than exact numbers, with red points indicating higher values and blue points indicating lower values for a specific particle’s feature. Grey points indicate that the values for that feature are not present, which happens for particles that were not within the search radius in the past snapshot. The radius of the particle at the current and past snapshot has the largest influence on the model’s classification for both populations, indicated by the larger SHAP values the model assigns this feature. The present radial and tangential velocities have similar shapes for both groups while past velocities only significantly contribute with extreme velocities.
Figure 7.— Comparison of kinetic-energy cuts to Sparta’s classification. We investigate cuts in linear (top-row) and log (bottom-row) scaled phase-space of $\ln(v^{2}/v^{2}_{\rm 200m})$ versus $r/R_{200m}$ from all the WMAP7 simulations. The particles are split by their classification according to Sparta and then further by their sign of their radial velocity. The two kinetic-energy cut methods, described in Section IV.2, are overplotted as green lines (Fast) and cyan bins (Optimized). The kinetic-energy cuts classify particles below their lines as orbiting and those above as infalling. The blue line indicates the furthest radial extent that particles are considered for the simple kinetic-energy cuts’ fit. The bin-by-bin fit of the Optimized method disagrees with the best-fit lines significantly, and naturally leads to a more accurate match to Sparta’s classification.
@@
Overall, the model seems to classify particles without past information as infalling, as one would expect for particles that were outside the search radius at that time. The only exceptions are particles that belong to halos that did not exist 1.0 dynamical time ago or had fast enough infalling velocities to have a pericenter despite the distance. Only a small subset of this population eventually orbits the halo, reinforcing our decision to limit the snapshot separation to 1.0 dynamical time. Increasing the time separation would increase the number of particles in this category, reduce the relevance of past snapshots, and provide less information on a particle’s orbiting classification.
IV.2. Comparison to Kinetic Energy Cut
@@
Figure  7 compares the Fast method (green lines) to the distribution from Sparta in isolated halos, defined as halos whose most massive neighbor within $2R_{\rm 200m}$ is at most 20% of the mass of the central halo. We split the particle population into four samples with positive/negative radial velocities and orbiting/infalling classifications from Sparta. The green lines in Fig. 7 show that the linear kinetic-energy cut separates the bulk of the orbiting and infalling populations as classified by Sparta. The figure shows both linear (top) and logarithmic (bottom) scales, highlighting the bulk of the population and outliers, respectively. Figure 8 additionally shows the fraction of orbiting particles as a function of radius as determined by both algorithms. The Fast and Sparta algorithms are in reasonably good agreement, except for the very inner regions and beyond $2R_{\rm 200m}$, where Fast classifies all particles as infalling.
@@
By contrast, the logarithmic scale for the bottom row of Fig. 7 highlights the tails of the distributions, making it easier to see where the Sparta and Fast classifications disagree. The most important difference is due to a population of particles classified as infalling by Sparta at large radii and with low kinetic energy, i.e the particles that fall below the Fast line in the right-most plot of Fig. 7. These particles “boost” the outskirts of the orbiting profile of halos in the Fast algorithm relative to the profile obtained with Sparta. This is illustrated in Fig. 9, where we compare the orbiting and infall profiles of halos as determined using Fast and Sparta. The truncation of the Sparta orbiting profiles is much more sensitive to peak height than the truncation observed using the Fast algorithm. This type of sensitivity is expected: the outer profile is known to be sensitive to accretion rate (diemer_14), which increases with peak height, suggesting the outer profiles in Sparta are better physically motivated than those obtained using the Fast algorithm.
However, it is not immediately clear whether this is a limitation of the Fast algorithm in particular or kinetic-energy cuts in general. We have thus also tested the Optimized method, which corresponds to finding a radius-dependent kinetic energy cut that minimizes the differences with Sparta’s classification. Figure  7 shows that the resulting cuts (in cyan) cannot be adequately described as a line. Naturally, the Optimized classification better matches Sparta, but the right columns of Fig. 9 demonstrate that differences near the truncation of the orbiting profiles remain. These differences are fundamentally due to the fact that distant, low-energy infalling particles have phase-space positions indistinguishable from orbiting particles near their apocenter, leading phase-space based classifiers to label these particles as orbiting. However, when adding information from a snapshot that is 1.0 dynamical time earlier, it becomes trivial to separate these particles: infalling particles used to be much further away from the halo, while orbiting particles in the halo outskirts used to be near the halo center at that time. We have explicitly verified this is the case, explaining why our machine-learning method can adequately distinguish between orbiting and infalling particles in the halo outskirts.
@@ Acknowledgments
This research was supported in part by the National Science Foundation under Grant numbers 2206690. BD is grateful to the Sloan Foundation for their financial support. E.S. and E.R. were supported by NSF grant 2206688. E.R. is also funded by DOE grant DE-SC0009913. The Erebos simulations were run on the Midway computing cluster provided by the University of Chicago Research Computing Center, and our analysis was performed on the Zaratan cluster at the University of Maryland. This research extensively used the python packages Numpy (code_numpy2), Scipy (code_scipy), Matplotlib (code_matplotlib), Pygadgetreader (Thompson, 2014), XGBoost (Chen and Guestrin, 2016), SHAP (Lundberg et al., 2019), and Colossus (diemer_18_colossus).
@@
Our code is publicly available at github.com/ZevVladimir/ATHENA or at 10.5281/zenodo.17228903 (catalog 10.5281/zenodo.17228903), including a description of how to create the datasets needed and how to train additional models. We also provide the json of the XGBoost model used and configuration parameters. The full particle data for the Erebos $N$-body simulations are too large to be permanently hosted online, but they are available upon request.